# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v1

@@ 2.3 Proof of the Convergence
(39) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}<\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the distance between the iterating points and the optimum is always decreasing, since the distance is always non-negative so that it has lower bound, $(\lambda_{k}(x),\mu_{k}(x))-(\lambda^{*}(x),\mu^{*}(x))$ is a convergent sequence so that $(\lambda_{k}(x),\mu_{k}(x))$ is bounded. According to Arzelà-Ascoli Theorem, there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, so that we cannot assert that the sequence is necessarily convergent in all cases and need careful discussion. One one hand, if there is only one optimum in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (38). If $\mathcal{G}$ has more than one optimum, then suppose there exists more than one sub-sequence, for example, $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$, $(\lambda_{k_{j}}(x),\mu_{k_{j}}(x))$ that converges to $(\lambda^{*}_{i}(x),\mu^{*}_{i}(x))$, $(\lambda^{*}_{j}(x),\mu^{*}_{j}(x))$ respectively, then due to (34) we show that although the sequence might have converged subsequence that approaches to different optimum, the adjacent term tends to zero so that the adjacent primal solution in iteration also tends to zero. On the other hand, as numerical algorithms will introduce error inevitably, if after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ is exactly the unique optimum so that the right hand of (39) is zero, then from monotonicity of iteration we know $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. Even if $\mathcal{G}$ has more than one optimum and after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ approaches one of the optima $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$, since there is another optimum $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
@@
thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ is always convergent to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
Now we turn to prove the convergence of series $g_{k}(x)$. As each $g_{k}(x)$ is obtained by solving an ODE function as the optimum with respect to a variational problem, we should first prove the existence and uniqueness to the solution of $g$ for each iteration.
@@
The solution of (11) with $g(0)=0,g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
Proof 2.6.
Note that the differential equation of the boundary value problem we are solving has only zero solutions under the corresponding homogeneous differential equation and homogeneous boundary conditions, in order to prove the existence and uniqueness of our solution in each iteration [3]. Recall that when dealing with the homogeneous condition, we exactly have
where
(43) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
It is obvious that $C_{3}$ and $C_{4}$ is zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
(44) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
If $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (44) hold. If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ is zero. As a consequence, during each iteration, there always exist a unique solution to $g(x)$, so that we can simply discuss the convergence based on the continuous-dependency between the solution and the coefficients.
Corollary 2.7.
For each iteration of (13), there always exists a unique solution.
Now we begin to prove the convergence of $g_{k}(x)$. For this property during the iteration in Algorithm 1, we have the following theorem:
Theorem 2.8.
Let $f(x)$ be first-ordered Lipschitz-continuous function that satisfies $|f^{\prime}(x)-f^{\prime}(y)|\leq L|x-y|$ for any $x,y$, and suppose $g_{0}(x)$ satisfies $0\leq g_{0}(x)\leq f(x)$. If $\{g_{k}(x)\}$ satisfies
@@
[…]
Then if the step $\eta$ is sufficiently small so that $\eta<\frac{1}{L}$ for each iteration, $g_{k}(x)$ must terminate at one of the optimum.
Proof 2.9.
Based on lemma 15, theorem 2.3 and theorem 2.5 we have shown that our original problem has unique solution during each iteration. As (11) is a linear ODE, its solution is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ converges or falls early in one of the optima, and
[…]
$\{(\lambda_{k}(x),\mu_{k}(x))\}$ has sub-sequences approaching different optimum respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
@@ 3 Implementation
3.1 Algorithm
In this part we state the whole implementation of our method. Our implementation can be mainly divided into three parts. Given a signal to be decomposed, we first use variational method in terms of iteratively constructing such a function sequence that each term of it is the optimal solution obtained by vanishing the variation of the Lagrangian Slack Function under the current evolutionary state with respect to inequality constraint conditions and then updating the weight of constraint using dual ascent method. Thirdly, we apply Kernel Density Estimation (KDE) to the amplitude values of each point in the spectrum to eliminate residual gaps at the bottom, which is based on the observation that the residual gap at the bottom will results in a clustering behaviour around the corresponding value that the residual gap share. The proper threshold to eliminate the gap for further evaluating the number of IMFs should locate where the value of the KDE function achieves the maximum. In short, our whole workflow can be summarized as in Algorithm 2.
@@ 3.1 Algorithm
3: Pre-process by extending $f(x)$ at both ends.
@@
3.2 Pre-processing
In order to solve the equation effectively, before we start the iteration, we first normalize the signal into $[0,1]$, in both domain and range. Since in each iteration, our optimal baseline should satisfy the boundary condition of the ODE, so that when dealing with high-pass or low-pass signals where spectrum has sharp slope at the ends, the to-be-evaluated baseline will also exhibits the sharp slope, which will in turn have bad effect on the convergence at the ends. To overcome the boundary effect, we extend our original function by making a gentle decay at both ends of the spectrum that not only completes the peaks at both ends but also alleviate the sharp slope effect. We will discuss this in detail later in our experiments.
3.3 Post-processing
Despite the fact that the iteration procedure to find the supporting baseline converges in reasonable time in most cases when applied with proper extension to the original spectrum of the signal, the shape of the spectrum do have great impact on the convergence rate. For example, since our optimization strategy discourage the curvature (via second-order derivative) of the curve, the optimal baseline is prone to converge more faster at points where original spectrum is flat. On the other side, although $g(x)$ will eventually adhere closely to the lower bound of $f(x)$, it is difficult to guarantee that the error between $f(x)$ and $g(x)$ calculated after stopping the iteration is consistent at every point so that f(x)g(x)f(x)-g(x) will have some non-zero points where it should be zero, which makes further retrieving modes of difficulty. Since $g(x)$ converges to the bottom of $f(x)$, it is reasonable to assume that the gap is typically small so that f(x)g(x)f(x)-g(x) is almost zero except for those frequency points corresponds with effective signal components. In order to eliminate the gap, we use Kernel Density Estimation(KDE) method to estimate the best threshold, so that the gap of f(x)g(x)f(x)-g(x) will be exposed where the value of the KDE function achieves the maximum. Fig. 2 illustrates such a phenomenon.
Figure 2: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function. Fig. 2a shows the converged $g(x)$ and Fig. 2b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 2c shows the estimate kernel density pertaining the gap at the bottom. 2d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
@@ 4 Numerical Results
In this section we first show our experiments to our method. In order to prove the effectiveness of our method, we choose some standard function that were repeatedly introduced as benchmarks in some pioneer works [5, 13, 12], yet pay more attention on investigating the evolution of our algorithm during the course of iteration, to show the convergence and robustness of our method. Since our algorithm can be reckoned as an algorithm for finding the number of intrinsic modes and initial center frequencies that the proceeding VMD module needs, our experimental focus is on exploring the behaviour of the algorithm towards various signals in the process of finding the baseline and talking about the KDE process for eliminating the gap, rather than elaborating the performance brought by the subsequent VMD decomposition in the final decomposition results. Then we turn to discuss why we should and how we eliminate the boundary effect by extending the spectrum obtained from the FFT of original signals.
@@ 4.1 Experiments on Finding the Supporting Baseline
In this part we show our experiments on finding the supporting baseline for some benchmark signals that were frequently used to test the performance of algorithms in several articles [5, 12]. As default, all signals are obtained by linearly normalizing the domain of the original function to between 0 and 1, and the range of which is normalized into 0 and 1 before iterations to keep the stability. We uniformly add 200 points to the domain and only preserve half of the transformed spectrum as our initial spectrum to be iterated according to the properties of Fourier Transform with respect to real signals. In this way, the Nyquist frequency is 100Hz, which means the highest signal angular frequency that can be effectively represented is $200\pi$, which corresponds to 1 in our normalized FFT results. Our convergence criterion is $\cfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. Due to the space limitations, we only illustrate the final convergent state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error, and indicate the number of iteration steps in the caption. The intermediate iteration process status and the decomposed signal is given in detail in the supplementary materials 5.
Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency. Here we choose $y(t)=100\sin 20\pi t$. As we see the normalized center frequency is $0.1$.
Figure 3: The initial, final convergent state and convergence curve in logarithmic coordinate system with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
@@
Figure 4: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
as our signal. It has a low frequency component, $6t^{2}$, two pure harmonic components, $\cos 60\pi t,\cos(80\pi t-10\pi)$, in different intervals that does not intersect with each other, respectively, and one narrow banded component, $\cos(10\pi t+10\pi t^{2})$, in the whole interval. To our surprise, the iteration converges even more quickly than those signals that has fewer modes. In our cases, it takes about 15000 iterations to achieve the convergence.
Figure 5: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to time series (45). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
Experiment 4: Now we show some examples on signals that conflict with the narrow-band assumption. We take $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. As we see that, the spectrum of the signal exhibits two main peaks, however the signal is not a completely narrow-band signal since it exhibits comb-shaped spectrum with respect to harmonic components, although the energy of those harmonic components are relatively very small(Fig. 6). Our algorithm extracts three components of the signal, despite that in the mainstream view there should be only two IMFs. However, since it is not a narrow-band signal, the difficulty of decomposition objectively exists since one can not even provide a specific standard to determine whether these small harmonic components should be considered as an intrinsic mode. We also observe that, even in [5, 12], when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it. Despite the fact that our results has three components, every separate component exhibits better purity than the pioneer works. On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours (0.1 ours vs 0.2 theirs, see S5.13a), which also indicates that our algorithm has better performance in this case.
Figure 6: The final convergent state in logarithmic coordinate system for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
@@
Figure 7: The final convergent state in logarithmic coordinate system for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
4.2 Experiments on Extension of Original Spectrum at Both Ends
Although supporting baseline can be retrieved by the method stated above, in real cases there are circumstances where convergence is very slow. For example, since the baseline must satisfy the boundary condition on both side, the trend of spectrum at both ends has a great impact to the shape of the iterating curve. In particular, for low or high pass signals of which the spectrum have a significant slope trend in low and high frequency regions, the iterating curve is prone to adhere the slope at both ends and continue to keep such inertia to some extent even when slipping out of those slope regions. This phenomenon brings us two bad effects in analysing signals: On one side, for low or high pass signals, their spectrum at both ends always exhibit half peaks since the scope of representable spectrum from FFT is always limited and the spectrum of those signals will inevitably hits the boundary, making the peak not complete and the estimation of center frequency inaccurate. On the other side, inertia effect will be harmful to the shape of the baseline, since in regions adjacent to low and high frequency bands, the function value of supporting baseline will have very large gap to the lower bound of the original spectrum due to the inertia of the downward trend. In order to circumvent such effect, we extend the original spectrum by forcing the spectrum to gently drop to the minimum of original spectrum and leave some extra flat intervals at both ends. This method not only take the advantage of trend effect, in that the curve will tend to keep horizontally at both ends which can make the peak be finally filtered out, but also complete the peaks at both ends without any harmful effects to the other regions as the experiment shows. Fig. 8 shows the evolution of spectrum baseline during the iteration with respect to the signal $y(t)=60t+60\cos 200\pi t$. This typical signal is a band-stop signal which shares both high and low pass feature. Note that if we don’t extend the spectrum, the baseline exhibits a sharp decreasing in the low frequency area and can not catch up the lower bound in the interval $[0.05,0.4]$ and the similar phenomenon also happens at the high-frequency area, even after 1500 iterations. This shows the necessity of proper extension to the original spectrum.
Figure 8: Iterations for finding the supporting baseline without extending the spectrum to the signal $y(t)=60t+60\cos 200\pi t$. The 500th(a), 1500th(b) iterations with no extension and the 500th(c), 1500th(d) iterations with extension are shown respectively. The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.

### v2

@@ 2.3.3 Convergence in Dual Space
(52) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the distance between the iterating points and the optimum is always decreasing, since the distance is always non-negative so that it has lower bound, $(\lambda_{k}(x),\mu_{k}(x))-(\lambda^{*}(x),\mu^{*}(x))$ is a convergent sequence so that $(\lambda_{k}(x),\mu_{k}(x))$ is bounded. According to Arzelà-Ascoli Theorem, we can only assert that there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, which does not indicate that the whole iterating sequence is necessarily convergent in all cases and needs careful discussion. There are four cases when the iterating sequence is approaching the optimum:
CASE I: If there is only one optima in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (51).
CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (51) and (45) we know that the value of $\mathcal{G}$ is approaching the minima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero. Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, although the classical convergent behaviour for the sequence in this case might not be guaranteed.
CASE III: If the iterating sequence has approached to the unique optima during some steps, then due to (52) the next iteration will still stay at the unique optima.
CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then since in the convex set there are at least another optima $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
@@
thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$.
In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ can always terminate at the point as close as possible to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
2.3.4 Convergence in Primal Space
@@
Proof 2.11.
As the solution of (10) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k}(x)$ is approaching $g^{*}(x)$ when $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
@@ 3 Implementation
In this chapter, we focus on the detail about the implementation and numerical scheme about the algorithm for evaluating problem (10). Our workflow consists of 4 interconnected stages: (1) We discretize (19) by using Finite Difference Method,[20] transforming it into a linear system. (2) Crucially, by expressing the derivations of both solution $g(x)$ and the weight function $\alpha(x)$ as linear combination of their nodal values, we recast the entire ODE into a compact matrix form by introducing component-wise multiplication with a mild extension of Hadamard production. By imposing well-posed boundary conditions, we make the coefficient matrices full-rank and the whole problem invertible so that the evaluation can be significantly accelerated by modern software toolkits via matrix operation. (3) To further accelerate the convergence and make algorithm effective when handling high or low-passed signals, a smooth extrapolation technique is applied to the input spectrum at the both ends. (4) Finally, after obtaining the optimal baseline, a statistically grounded threshold determined via kernel density estimation (KDE), is employed to automatically identify significant spectral peaks, thereby completing the mode extraction. The following subsections detail each of these components.
3.1 Spatial Discretization Scheme
We discretize the solution domain $\Omega=[a,b]$ into $N+1$ equally spaced nodes $x_{0},x_{1},\dots,x_{n}$, where $x_{0}=a$ and $x_{n}=b$, with the grid spacing $h=\dfrac{b-a}{N}$, thus the value of any continuous function $p(x)$ at the nodes are recorded as a column vector $\boldsymbol{p}=[p(x_{0}),p(x_{1}),\dots,p(x_{N})]^{\top}\in\mathbb{R}^{N+1}$. Then we employ Five-Point Central Difference Scheme to transform different orders of derivatives into the linear combination with respect to the primal nodal values. For instance, for the fourth-order derivative $g^{(4)}(x)$ at interior nodes $x_{i},i=2,3,\dots,N-2$ we have
(55) $g^{(4)}(x_{i})\approx\cfrac{g(x_{i+2})-4g(x_{i+1})+6g(x_{i})-4g(x_{i-1})+g(x_{i-2})}{h^{4}}$
and the corresponding boundary conditions can be approximately expressed as
(56) $\left\{\begin{aligned} g(0)&=g(x_{0})\\ g(n)&=g(x_{N})\\ g^{\prime}(0)&\approx\cfrac{g(x_{1})-g(x_{0})}{h}\\ g^{\prime}(n)&\approx\cfrac{g(x_{N})-g(x_{N-1})}{h}\end{aligned}\right$.
Note that (55) and (56) can be transformed into compact matrix form via a pentadiagonal matrix $\boldsymbol{G}^{(4)}\in\mathbb{R}^{(N+1)\times(N+1)}$, thus we have
(57) $\boldsymbol{G}^{(4)}=\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\end{bmatrix}$
thus $\boldsymbol{G}^{(4)}\boldsymbol{g}$ is the fourth-ordered derivative of $\boldsymbol{g}$ at each node and the non-zero entry stands for the corresponding coefficients of difference and boundary conditions. $\boldsymbol{G}^{(3)}$, $\boldsymbol{G}^{(2)}$, $\boldsymbol{G}^{(1)}$ and $\boldsymbol{A}^{(4)}$ to $\boldsymbol{A}^{(1)}$ have the similar structure as $\boldsymbol{G}^{(4)}$, we give the detail of each pentadiagonal matrix in A.
By integrating both the difference strategy and the boundary condition (19) is transformed into (21), where $\odot$ stands for component-wise multiplication. However, in (21) $\boldsymbol{g}$ is inseparable so that we cannot resort to effective evaluation. To solve this, our $\odot$ is just an extension of standard Hadamard production by adding some compatible broadcasting rules in dealing with column vector, that is $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
(58) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$. By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$, we can transform (21) into
(59) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})\boldsymbol{g}+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})\boldsymbol{g}+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})\boldsymbol{g}=\boldsymbol{\beta}-\boldsymbol{\lambda}+\boldsymbol{\mu}$
so that we can treat (21) as common matrix equation and get the solution by simply using matrix inversion with modern mathematical software tookits.
Figure 2: Iterations for finding the supporting baseline without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum. The solid line is the original spectrum and the dotted line is the evaluated supporting baseline. With smooth extrapolation to the original spectrum, the speed of convergence has significant improvement.
3.2 Treatment of Boundary Conditions
The primal problem requires the supporting baseline $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$. For low and high pass signals of which the spectrum value are high and drop abruptly at either of the ends, this will make $g(x)$ inherits the trend of $f(x)$ at both ends at some interval so that convergence in this case will be very slow with reasonable grid space. To alleviate this effect, we adopt a smooth extrapolation strategy by symmetrically appending gently decaying segments at both ends to make the extended value align with the minimum of $f(x)$, and the extended derivative smoothly approach to zero within the extended region. This strategy significantly improves the condition number of the coefficient matrix of the discrete system, ensuring the numerical stability of the iterative solution, without any altering the primal spectrum within the original interval. Fig. 2 shows the illustrations.
3.3 Postprocessing: From Baseline to Mode Identification
Upon the above course when the algorithm converges, we obtain the optimal discrete baseline vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly. However, there are some residual background fluctuations at the bottom that hinders one to objectively distinguish those modes. To solve this problem, we assume that the residual consists of two components: first, a background fluttering component originate from noise or minor undulations, of which the amplitude concentrates within a very small range; second, significant protrusions from original modes, which are larger compared with the former and more scattered. In this view, when investigating the value of the residual spectrum value distribution, we can intuitively find that most values concentrate near the level where fluttering component resides. This inspires us to employ Kernel Density Estimation (KDE) to non-parametrically model the spectrum value distribution as a probability density function:
(60) $\phi(t)=\cfrac{1}{mh}\sum\limits_{i=1}^{m}K\left(\cfrac{t-r_{j}}{h}\right)$
where $K(\cdot)$ is Gaussian function, with $h$ the bandwidth and $r_{j}$ the discrete nodes in the residual vector $\boldsymbol{r}$, and we evaluate the best threshold to eliminate the fluttering noise as
(61) $t^{*}=\underset{t}{\arg\max}\phi(t)$
so that we can isolate each mode in the function $f(x)-g^{*}(x)-t^{*}$.
Figure 3: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function. Fig. 3a shows the converged $g(x)$ and Fig. 3b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 3c shows the estimate kernel density pertaining the gap at the bottom. 3d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
3.4 The Whole Workflow
Based on the above discrete framework, each step of Algorithm 1 can be implemented by specific linear algebraic operations. The whole workflow can be summarized as Algorithm 2. We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the supporting baseline function $g(x)$. After that we smoothly extend $f(x)$ at both ends and begin Algorithm 1 until it converges. Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, so as to isolate band components.
@@ 3.4 The Whole Workflow
3: Smoothly Extending $f(x)$ at both ends.
@@ 4 Numerical Results
In this section, we show some experiments to validate the applicability of our methods. Our experiments focus on 3 goals: (1) To validate the convergence and robustness of our method under synthetic signals like single-modality signal, dual-modality signal, segmented signal, comb-spectrum signal, dense-modality signal to demonstrate the extensive adaptability of the model across modes with various distribution. (2) To compare our method against the most challenge-able method Successive Variational Mode Decomposition (SVMD), another recursive, adaptive method determining the intrinsic modes, exhibiting the advantage of our method on not generating redundant modes or eliminating necessary modes. (3) To understand the applicability of our method in real-world signals like electrocardiogram (ECG) signals, showing the accuracy of decomposition by investigation of the signals reconstructed from the decomposed modes.
All experiments are conducted under the same parameter settings unless otherwise stated. The stopping criterion is set as $\dfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$. The number of discrete grid points is fixed at $N=200$ for the normalized frequency domain $[0,1]$. Detailed descriptions of the test signals and comparative protocols are provided in the corresponding subsections.
Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
@@ 4.1 Experiments on Finding the Supporting Baseline
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. We only illustrate the initial and final state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in C.
Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency. Here we choose $y(t)=100\sin 20\pi t$. As we see the normalized center frequency is $0.1$. The algorithm takes 5161 steps and 1.88 seconds to converge.
@@
Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
as our signal. It has a low frequency component, $6t^{2}$, two pure harmonic components, $\cos 60\pi t,\cos(80\pi t-10\pi)$, in different intervals that does not intersect with each other, respectively, and one narrow banded component, $\cos(10\pi t+10\pi t^{2})$, in the whole interval. To our surprise, the iteration converges even more quickly than those signals that has fewer modes. In our cases, it takes about 3026 iterations to achieve the convergence.
Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to time series (62). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
Experiment 4: Now we show some examples on signals that conflict with the narrow-band assumption. We take $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. As we see that, the spectrum of the signal exhibits two main peaks, however the signal is not a completely narrow-band signal since it exhibits comb-shaped spectrum with respect to harmonic components, although the energy of those harmonic components are relatively very small(Fig. 7). Our algorithm extracts three components of the signal, despite that in the mainstream view there should be only two IMFs. However, since it is not a narrow-band signal, the difficulty of decomposition objectively exists since one can not even provide a specific standard to determine whether these small harmonic components should be considered as an intrinsic mode. We also observe that, even in [5, 12] when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it. Despite the fact that our results has three components, every separate component exhibits better purity than the pioneer works. On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours, which also indicates that our algorithm has better performance in this case.
Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
Figure 7: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual. As compared with (b) and (c), for original VMD, both the results has reconstruction error twice as large as ours.
@@
Figure 8: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.